please solve the system of linear equation by elimination method
x-y/x+y=4/5
7x-13y=50

i need its answer today..please

Found 2 solutions by best82, Alan3354:
Answer by best82(13)

You cannot eliminate this.

Answer by Alan3354(69443)

please solve the system of linear equation by elimination method
x-y/x+y=4/5
7x-13y=50
------------
x-y/x+y=4/5
cross multiply
5(x-y) = 4(x+y)
5x-5y = 4x+4y
x - 9y = 0
Times 7 -->
7x - 63y = 0
7x - 13y = 50
-------------- Subtract
0x - 50y = -50
y = 1
=================
x - 9y = 0
x = 9